Shift report 20000808 00:00-08:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Tue Aug 08 2000 - 07:58:22 EDT

  • Next message: Peter H. L. Christiansen: "Copying files to home disks"

    ------------------------------------------------------------
    Date: Tue Aug  8 07:58:22 2000, Shift: 00:00-08:00
    Supervisor: DO, with: ZIS
    Log book#4 pages 29 - 29
    
    Shift Summary:
    ------------------------------------------------------------
    1. MC worked on injection during all night. No useful beam
       has been obtained. A couple of ramps were unsuccessful.
       Ramiro Debbe called around 1:00 to tell us to turn off
       D5 because of the 2 tiles previously mounted for 
       calibration.
       Finally, at 7:40, beam was successfully ramped, stored 
       and cogged. Gamma = 70, Intensities: B = 806, Y = 170. 
        
    2. Run 2106 started at 7:41
    3. Noticed an oddity in BRAHVO that occured only once:
       wanted to check H1 and C1 but connection to mainframe
       refused. Closed BRAHVO and restarted it -> OK. 
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Tue Aug 08 2000 - 08:01:38 EDT